# The fundamental localization phases in quasiperiodic systems: A unified framework and exact results

### v2

@@ Abstract
The disordered quantum systems host three types of quantum states, the extended, localized, and critical, which bring up various distinct fundamental phases, including the pure phases and coexisting ones with mobility edges. The quantum phases involving critical states are of particular importance, but are less understood compared with the other ones, and the different phases have been separately studied in different quasiperiodic models. Here we propose a unified framework based on a spinful quasiperiodic system which unifies the realizations of all the fundamental Anderson phases, with the exact and universal results being obtained for these distinct phases. Through the duality transformation and renormalization group method, we show that the pure phases are obtained when the (emergent) chiral symmetry preserves in the proposed spin-1/2 quasiperiodic model, which provides a criterion for the emergence of the pure phases or the coexisting ones with mobility edges. Further, we uncover a new universal mechanism for the critical states that the emergence of such states is protected by the generalized incommensurate matrix element zeros in the spinful quasiperiodic model, as a novel generalization of the quasiperiodic hopping zeros in the spinless systems. We also show with the Avila’s global theory the criteria of exact solvability for the present unified quasiperiodic system, with which we identify several new quasiperiodic models derived from the spinful system hosting exactly solvable Anderson phases. In particular, we reach a single model that hosts all the seven fundamental phases of Anderson localization. Finally, an experimental scheme is proposed to realize these models using quasiperiodic optical Raman lattices. This study establishes a unified theoretical framework which enables an in-depth exploration of the broad classes of all fundamental localization phases in quasiperiodic systems, and offers key insights for constructing their exactly solvable models with experimental feasibility.
@@ I Introduction
Anderson localization, wherein quantum states localize exponentially in real space due to disorder-induced scattering, is a fundamental and universal phenomenon in disordered systems (Anderson, 1958; Lee and Ramakrishnan, 1985; Kramer and MacKinnon, 1993; Evers and Mirlin, 2008). It manifests distinctly across two categories of aperiodic systems: disordered systems and quasiperiodic lattices. In disordered systems, scaling theory predicts that noninteracting states are localized in one and two dimensions even under weak disorder, whereas in three dimensions, an Anderson transition from extended to localized states occurs only with sufficiently strong disorder (Abrahams et al., 1979). In contrast, quasiperiodic (QP) lattices exhibit richer physics including Anderson transitions in all spatial dimensions, governed by quasiperiodic parameters (Aubry and André, 1980; Roati et al., 2008). Importantly, both extended and localized states can coexist within a single quantum phase, separated by characteristic energies known as mobility edges (MEs) (Soukoulis and Economou, 1982; Das Sarma et al., 1988; Boers et al., 2007; Biddle and Das Sarma, 2010; Ganeshan et al., 2015; Li et al., 2017; Gopalakrishnan, 2017; An et al., 2018; Lüschen et al., 2018; Kohlert et al., 2019; Deng et al., 2019; Wang et al., 2020a; Li and Das Sarma, 2020; Duthie et al., 2021; An et al., 2021; Roy et al., 2021; Roy and Sharma, 2021; Wang et al., 2021a, 2022a; Longhi, 2023; Wang et al., 2023; Gao et al., 2023; Longhi, 2024a, b; Wei et al., 2024; Gao et al., 2024; Liu et al., 2024; Tabanelli et al., 2024; Shimasaki et al., 2024a). The ability to derive exact analytical expressions for characteristic length sclaes and MEs in various QP models Wang et al. (2020a, 2021a); Zhou et al. (2023); Gonçalves et al. (2023a); Liu et al. (2022) plays a crucial role in advancing the fundamental understanding of localization transitions, providing theoretical insights that extend beyond numerical simulation.
@@
Between the extended and localized states lie the critical states, which are delocalized in both position and momentum spaces. Critical states have recently attracted extensive research interests (Hatsugai and Kohmoto, 1990; Han et al., 1994; Tanese et al., 2014; Zhou et al., 2023; Wang et al., 2016; Liu et al., 2015; Zeng et al., 2016; Yao et al., 2019; Wang et al., 2020b; Goblot et al., 2020; Xiao et al., 2021; Liu et al., 2022; Li et al., 2023; Wang et al., 2022b; Gonçalves et al., 2023a; Lin et al., 2023; Lee et al., 2023; Dai et al., 2023; Shimasaki et al., 2024b; Li et al., 2024; Wang and Zhou, 2025; Huang et al., 2025; Chen et al., 2024; Guo, 2024; Dotti et al., 2024; Yang et al., 2024; Duncan, 2024; Yao et al., 2024; Bai and Weld, 2025) due to their unique properties, such as local scale invariance, multifractality, and critical quantum dynamics. Recent advances, particularly through Avila’s global theory (Avila, 2015; Zhou et al., 2023), refine a rigorous characterization Simon and Spencer (1989); Jitomirskaya and Marx (2012) of the critical states, showing that critical states generally arise when the hopping couplings in a QP system are non-uniform and possess incommensurately distributed zeros (IDZs) in the thermodynamic limit Liu (2024). The exact theory enables unambiguous determination of critical phases and the new MEs between critical and other states Zhou et al. (2023). Furthermore, this rigorous characterization of critical states provides a well-defined foundation for incorporating many-body interactions, thereby extending the paradigm of criticality beyond the single-particle picture. Inclusion of interactions further enriches the physics of the critical phase, with multifractality of the wave function influencing both ground state properties associated with symmetry breaking (Liu, 2024; Feigel’man et al., 2007; Burmistrov et al., 2012; Zhao et al., 2019; Sacépé et al., 2020; Gonçalves et al., 2024) and the emergence of many-body critical (MBC) phases (Wang et al., 2021b, 2020b) at infinite temperature. The MBC phase, interpolating the thermal phase and many-body localization (MBL) (Pal and Huse, 2010; Nandkishore and Huse, 2015; Schreiber et al., 2015; Bordia et al., 2016, 2017a; Lüschen et al., 2017; Bordia et al., 2017b), signifies a non-ergodic but delocalized regime that violates the eigenstate thermalization hypothesis (ETH) (D’Alessio et al., 2016; Deutsch, 1991; Rigol et al., 2008; Srednicki, 1994).
@@
The recently developed quasiperiodic mosaic lattice models (Wang et al., 2020a; Zhou et al., 2023; Gao et al., 2024; Huang et al., 2025) and their generalizations Longhi (2023); Dai et al. (2023); Lin and Gong (2024); Guo (2024); Wei et al. (2024); Longhi (2024a) advance the rigorous characterization of all eigenstates in the spectra through the analytical study of their localization properties (Lyapunov exponent or LE) using Avila’s global theory (Avila, 2015). With the mosaic models, the MEs can also be analytically obtained. On the other hand, while the various exactly solvable models are proposed, these different models are treated on a case-by-case basis. There is no unified theory showing that under what conditions a generic exactly solvable model undergoes phase transitions between the pure localized, extended, and critical phases without MEs, and otherwise it possesses MEs that separate these states, i.e., the transitions are energy-dependent. Besides, Avila’s global theory enables an analytic study of the spinless models with nearest-neighbour hopping, while the exact results for more complicated systems are scarce, hindering the in-depth understanding of the localization physics in more generic systems. For instance, the universal mechanism for critical states in the complex systems with internal degrees of freedom, such as spins, remains unknown. Addressing these issues is not only important in fundamental theory, but also of high significance in experiment, as the remarkable experimental progresses in engineering the spin and orbital degrees of freedom Zhang and Liu (2018); Liu et al. (2014); Wu et al. (2016); Wang et al. (2018); Sun et al. (2018); Song et al. (2019); Lu et al. (2020); Wang et al. (2021c); Zhang et al. (2024) have enabled the realization of complex QP systems beyond toy models. These crucial fundamental issues motivate us to establish a universal theory that can generically characterize spinful QP lattice systems without relying on specific microscopic details.
@@
Figure 1: Illustration of generic framework. (a) The different processes of the spinful quasiperiodic systems. The hopping coupling matrix $\Pi_{j}^{s,s^{\prime}}$ denote the momentum transfer coupling with the diagonal (off-diagonal) terms for the same (different) internal degrees of freedom. The on-site matrix $M_{j}^{s,s^{\prime}}$ represents the on-site spin flipped and on-site modulation of the system. (b) The conditions when the Avila’s global theory can give rise to the analytical characterization of the (reduced) 1D quasiperiodic chain. The nonzero acceleration of the complexified Lyapunov exponent $\gamma_{\epsilon}$ is quantized to the unique integer: $d\gamma_{\epsilon}/d\epsilon=2\pi\mathbb{Z}$. Then the original Lyapunov exponent can be obtained by the expression obtained from the $\epsilon\rightarrow\infty$.
@@ I.1 Summary of results
In this work, we propose to investigate the generic spin-$1/2$ quasiperiodic (QP) system, with which we establish a unified theoretical framework for all fundamental localization phases based on the exact and universal results uncovered in this system. The proposed generic QP framework not only unifies existing 1D spinful QP models with exact analytical solutions, but also encompasses spinless QP models through an additional Majorana representation. More importantly, this system establishes a versatile platform for constructing new exactly solvable QP models exhibiting rich phase structures, including pure phases and coexisting phases with MEs.
The universal results are obtained in this unified framework and can be summarized as follows. First, we show with duality transformation and renormalization group method that the pure extended, localized, or critical phases are obtained when the system preserves a chiral symmetry. Accordingly, the emergence of coexisting phases with MEs necessitates to break the chiral symmetry. Second, we unveil a new universal mechanism for critical states in the spinful quasiperiodic system, which arise from generalized incommensurate zeros (GIZs) in matrix elements. This mechanism is a novel generalization of the incommensurate hopping zeros in the spinless QP models, enabling a much broader pathway of realizing and characterizing the critical states. Third, we establish the exactly solvable condition, under which the 1D spinful QP system can be transformed into effectively spinless QP models of dressed particles, and then can be exactly solved. This result provides a powerful guiding principle to construct exactly solvable models that host a variety of novel localization phases with or without MEs, particularly the models encompassing all seven fundamental phases in localization physics.
@@ II.1 The spinful quasiperiodic model
$H=\sum_{j,s,s^{\prime}}\big{(}c_{j+1,s}^{\dagger}\Pi_{j}^{s.s^{\prime}}c_{j,s^{\prime}}+\mathrm{h.c.}\big{)}+\sum_{j,s,s^{\prime}}c_{j,s}^{\dagger}M_{j}^{s,s^{\prime}}c_{j,s^{\prime}},$ (1)
where $c_{j,s}^{\dagger}(c_{j,s})$ creates (annihilates) a particle on site $j$ with (pseudo)spin $s=\uparrow,\downarrow$ that can denote real spin, sublattices or particle-hole. This model includes two key ingredients, as illustrated in Fig. 1(a). First, the hopping coupling matrix $\Pi_{j}=\sum_{l}p_{j}^{l}\sigma_{l}$, as written in terms of Pauli matrices $\sigma_{l}$ with $l=\{0,x,y,z\}$, represents the spin-independent (for $l=0$) and the spin-dependent (for $l=x,y,z$) hopping coupling terms. The latter includes both the spin-conserved ($l=z$) and spin-flip ($l=x,y$) hopping terms, corresponding to the spin-orbit coupling in the periodic systems. Second, the onsite matrix $M_{j}=\sum_{l}m_{j}^{l}\sigma_{l}$ also involves the spin-independent chemical potential and the Zeeman coupling terms. The coefficients $p_{j}^{l}=t^{l}+\mu^{l}V_{j}^{\mathrm{od}}$ and $m_{j}^{l}=\lambda^{l}+V^{l}V_{j}^{\mathrm{d}}$, where $(t^{l},\lambda^{l},\mu^{l},V^{l})$ are constants, the off-diagonal quasiperiodic modulation $V_{j}^{\mathrm{od}}=\cos[2\pi\alpha(j+1/2)+k_{y}]$ and the diagonal quasiperiodic modulation $V_{j}^{\mathrm{d}}=\cos(2\pi\alpha j+k_{y})$. Here $\alpha=\lim_{n\rightarrow\infty}(F_{n-1}/F_{n})=(\sqrt{5}-1)/2$ is an irrational number approached by the Fibonacci sequence $F_{n}$, and $k_{y}$ is a constant phase factor. In this work, we take that the quasiperiodic term has a single frequency. To facilitate our discussions, we define the combinations of the Pauli matrices $\sigma_{l}$ as $\sigma_{\pm}=(\sigma_{x}\pm+i\sigma_{y})/2$ and $\Lambda_{\pm}=(\sigma_{z}\pm\sigma_{0})/2$.
This spinful QP system unifies various QP models, including the existing 1D spinful models with analytical solutions, and spinless models by taking the Majorana representation (see the details in Appendix C and Appendix D). More importantly, with this unified framework we shall develop a series of new and exactly solvable models which host various novel quantum phases with or without mobility edges. Below, we outline the three main theoretical approaches employed in this study.
@@ II.2 Dual transformation
The dual transformation is a non-local operation similar to the Fourier transform, and maps the system between real space and dual momentum space. It is defined as $c_{j,s}=(1/\sqrt{L})\sum_{n}e^{-i2\pi n\alpha j}c_{n,s}$. The extended, localized, and critical states behave differently under this transformation. The critical states exhibit behavior distinct from the other two. For an extended state, the wave function is delocalized in real space, but localized in dual momentum space, whereas the localized state is delocalized in dual momentum space but localized in real space. These opposing features are exchanged under the dual transformation. However, critical states remain delocalized in both real and dual spaces and are invariant under the transformation. For example, when the system has a QP hopping term $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$, which dominates over other hopping couplings and is invariant under the dual transformation due to the presence of incommensurately distributed zeros (IDZs) in the QP hopping modulation $V_{j}^{\rm od}$, the critical states may emerge and exhibit self-dual for having incommensurate nodes corresponding to the IDZs Zhou et al. (2023). These incommensurate nodes in the delocalized wave functions remain unchanged under the dual transformation, driving the delocalized states into critical states. An intuitive picture suggests that IDZs partition the 1D system into multiple segments, where the wave function reorganizes, exhibiting a self-similar pattern characteristic of critical states.
@@
The distinct behaviors of the three types of states can be mapped onto a 2D system, providing an intuitive physical interpretation, where $x$- and $y$-axes correspond to real and dual space, and dual transfomration coppresonds to an exchange between two axes. In this mapping, the extended state is delocalized along $x$-direction and localized along $y$-direction, and the opposite is true for localized states. The critical state is delocalized in both directions. Thus, the swap of the two axes transforms the localized and extended states, while leaving the critical states unchanged. We use 1D spinless 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) as an example, with the nearest neighbor hopping $t_{j}=t+\mu V_{j}^{\rm od}$ and on-site potential $V_{j}=V_{0}V_{j}^{\rm d}$ (see details in Appendix D). By interpreting the phase shift $k_{y}$ as the quasi-momentum in the $y$-direction and performing a Fourier transform, the 1D spinless EAA model maps to a 2D Hofstadter model that describes a particle moving in a magnetic field (Kraus and Zilberberg, 2012; Borgnia and Slager, 2023). In this mapping, the uniform hopping terms correspond to hopping in the $x$-direction, while the on-site potential $V_{j}^{\mathrm{d}}a_{j}^{\dagger}a_{j}$ corresponds to the hopping in the $y$-direction with a magnetic flux $\exp(i2\pi\alpha x)a_{x,y}^{\dagger}a_{x,y+1}+{\rm h.c.}$. Similarly, the QP hopping terms $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$ represent diagonal hopping in the 2D plane under the magnetic flux. The relative hopping amplitude shapes the cyclotron orbits while maintaining the flux within each orbit. Dominant hopping in the $x$- and $y$- directions in 2D facilitates the emergence of extended, localized states. In contrast, dominant diagonal hopping, giving equal weight to both directions, results in critical states.

### v4

@@ Abstract
The disordered quantum systems host three types of quantum states, the extended, localized, and critical, which bring up seven distinct fundamental phases in nature: three pure phases and four coexisting ones with mobility edges, yet a unified theory with full characterization and realization of all these phases has not been developed. Here we propose a complete and unified framework based on a spinful quasiperiodic (QP) system which realizes all the fundamental localization phases, with the exact and universal results being obtained for their characterization. First, we show that the pure phases are obtained when the chiral symmetry preserves in the proposed spinful QP model, giving a criterion for the emergence of the pure phases and otherwise the coexisting ones. Further, we uncover a novel universal mechanism for the critical states that their emergence is protected by the generalized incommensurate matrix element zeros in the spinful QP model, which considerably broadens the rigorous realizations of the exotic critical states. We then show the criteria of exact solvability for the present spinful QP system, with which we construct various exactly solvable models for all distinct localization phases. In particular, we propose two novel models, dubbed spin-selective QP lattice model and QP optical Raman lattice model, to achieve all basic types of mobility edges and all the seven fundamental phases of Anderson localization physics, respectively. The experimental scheme is proposed and studied in detail to realize these models with high feasibility. This study establishes a complete and profound theoretical framework which enables an in-depth exploration of the broad classes of all fundamental localization phenomena in QP systems, and offers key insights for constructing their exactly solvable models with experimental feasibility.
@@ I Introduction
Anderson localization, wherein quantum states localize exponentially in real space due to disorder-induced scattering, is a fundamental and universal phenomenon in disordered systems (Anderson, 1958; Lee and Ramakrishnan, 1985; Kramer and MacKinnon, 1993; Evers and Mirlin, 2008). It manifests distinctly across two categories of aperiodic systems: disordered systems and quasiperiodic lattices. In disordered systems, scaling theory predicts that noninteracting states are localized in one and two dimensions even under weak disorder, whereas in three dimensions, an Anderson transition from extended to localized states occurs only with sufficiently strong disorder (Abrahams et al., 1979). In contrast, quasiperiodic (QP) lattices exhibit richer physics including Anderson transitions in all spatial dimensions, governed by quasiperiodic parameters (Aubry and André, 1980; Roati et al., 2008). Importantly, both extended and localized states can coexist within a single quantum phase, separated by characteristic energies known as mobility edges (MEs) (Soukoulis and Economou, 1982; Das Sarma et al., 1988; Boers et al., 2007; Biddle and Das Sarma, 2010; Ganeshan et al., 2015; Li et al., 2017; Gopalakrishnan, 2017; An et al., 2018; Lüschen et al., 2018; Kohlert et al., 2019; Deng et al., 2019; Wang et al., 2020a; Li and Das Sarma, 2020; Duthie et al., 2021; An et al., 2021; Roy et al., 2021; Roy and Sharma, 2021; Wang et al., 2021a, 2022a; Longhi, 2023; Wang et al., 2023; Gao et al., 2023; Longhi, 2024a, b; Wei et al., 2024; Gao et al., 2024; Liu et al., 2024; Tabanelli et al., 2024; Shimasaki et al., 2024a). The ability to derive exact analytical expressions for characteristic length scales and MEs in various QP models Wang et al. (2020a, 2021a); Zhou et al. (2023); Gonçalves et al. (2023a); Liu et al. (2022) plays a crucial role in advancing the fundamental understanding of localization transitions, providing theoretical insights that extend beyond numerical simulation.
@@
Between the extended and localized states lie the critical states, which are delocalized in both position and momentum spaces. Critical states have recently attracted extensive research interests (Hatsugai and Kohmoto, 1990; Han et al., 1994; Tanese et al., 2014; Zhou et al., 2023; Wang et al., 2016; Liu et al., 2015; Zeng et al., 2016; Yao et al., 2019; Wang et al., 2020b; Goblot et al., 2020; Tong et al., 2021; Xiao et al., 2021; Liu et al., 2022; Li et al., 2023; Wang et al., 2022b; Gonçalves et al., 2023a; Lin et al., 2023; Yang et al., 2017; Lee et al., 2023; Dai et al., 2023; Shimasaki et al., 2024b; Li et al., 2024; Wang and Zhou, 2025; Huang et al., 2025; Chen et al., 2024; Guo, 2024; Dotti et al., 2024; Yang et al., 2024; Duncan, 2024; Yao et al., 2024; Bai and Weld, 2025) due to their unique properties, such as local scale invariance, multifractality, and critical quantum dynamics. Recent advances, particularly through Avila’s global theory (Avila, 2015; Zhou et al., 2023), refine a rigorous characterization Simon and Spencer (1989); Jitomirskaya and Marx (2012) of the critical states, showing that critical states generally arise when the hopping couplings in a QP system are non-uniform and possess incommensurately distributed zeros (IDZs) in the thermodynamic limit Liu (2024). The exact theory enables unambiguous determination of critical phases and the new MEs between critical and other states Zhou et al. (2023). Furthermore, this rigorous characterization of critical states provides a well-defined foundation for incorporating many-body interactions, thereby extending the paradigm of criticality beyond the single-particle picture. Inclusion of interactions further enriches the physics of the critical phase, with multifractality of the wave function influencing both ground state properties associated with symmetry breaking (Liu, 2024; Feigel’man et al., 2007; Burmistrov et al., 2012; Zhao et al., 2019; Sacépé et al., 2020; Gonçalves et al., 2024) and the emergence of many-body critical (MBC) phases (Wang et al., 2021b, 2020b; Yoshida et al., 2024; Ahmed and Roy, 2025; Roy et al., 2025) at infinite temperature. The MBC phase, interpolating the thermal phase and many-body localization (MBL) (Pal and Huse, 2010; Nandkishore and Huse, 2015; Schreiber et al., 2015; Bordia et al., 2016, 2017a; Lüschen et al., 2017; Bordia et al., 2017b), signifies a non-ergodic but delocalized regime that violates the eigenstate thermalization hypothesis (ETH) (D’Alessio et al., 2016; Deutsch, 1991; Rigol et al., 2008; Srednicki, 1994).
@@
The recently developed quasiperiodic mosaic models (Wang et al., 2020a; Zhou et al., 2023; Gao et al., 2024; Huang et al., 2025) and their generalizations Longhi (2023); Dai et al. (2023); Lin and Gong (2024); Guo (2024); Wei et al. (2024); Longhi (2024a) advance the rigorous study of localization physics in QP systems, but many important issues remain unexplored. The extended, localized, and critical states bring up seven fundamental localization phases in nature, including three pure phases and four coexisting ones (three doubly and one triply coexisting phases). However, thus far there was no unified quantum system which is shown to host all these phases. Especially, the exactly solvable models for the complete set of such phases are crucial for the analytical study of the generic localization physics, but are currently lacking. Moreover, there is no unified theory which exactly characterizes the phase transitions between the pure localized, extended, and critical phases, and all types of (conventional and anomalous) MEs. Besides, while Avila’s global theory enables an analytic study of the spinless models, generalization of the study to the cases beyond spinless systems is not only important in fundamental physics, but also desired in experiment. For instance, the universal mechanism for critical states in the complex systems with internal degrees of freedom, such as spins, remains unknown. On the other hand, the remarkable experimental progresses in engineering the spin and orbital degrees of freedom Zhang and Liu (2018); Liu et al. (2014); Wu et al. (2016); Wang et al. (2018); Sun et al. (2018); Song et al. (2019); Lu et al. (2020); Wang et al. (2021c); Zhang et al. (2024) have enabled the realization of complex QP systems beyond spinless toy models. These crucial fundamental issues motivate us to establish a universal theory that can generically characterize spinful QP lattice systems without relying on specific microscopic details.
@@
Figure 1: Illustration of generic framework. (a) The different processes of the spinful quasiperiodic systems. The hopping coupling matrix $\Pi_{j}^{s,s^{\prime}}$ denote the momentum transfer coupling with the diagonal (off-diagonal) terms for the same (different) internal degrees of freedom. The on-site matrix $M_{j}^{s,s^{\prime}}$ represents the on-site spin flipped and on-site modulation of the system. (b) The conditions under which Avila’s global theory yields analytical characterization of the (reduced) 1D quasiperiodic chain. The nonzero part of derivative of the complexified Lyapunov exponent $\gamma_{\epsilon}$ is quantized to the unique integer: $d\gamma_{\epsilon}/d\epsilon=2\pi\mathbb{Z}$. Then the original Lyapunov exponent can be obtained by the expression obtained from the $\epsilon\rightarrow\infty$.
@@ I.1 Summary of results
In this work, we propose the generic spin-$1/2$ QP system, with which we establish a unified theoretical framework for all fundamental phases in Anderson localization physics based on the exact and universal results obtained in this system. The proposed generic QP framework unifies not only the existing 1D spinful QP models with exact analytical solutions, but also the spinless QP models through an additional Majorana representation. Most importantly, this unified system establishes a versatile platform for constructing new exactly solvable QP models exhibiting rich phase structures, including all pure phases and all coexisting phases with MEs.
The universal results are obtained in this unified framework and can be summarized as follows. First, we show with duality transformation and renormalization group method that the pure extended, localized, or critical phases are obtained when the system preserves a chiral symmetry. Accordingly, the emergence of coexisting phases with MEs can be obtained by breaking the chiral symmetry. Second, we uncover a generalized universal mechanism for critical states in the spinful quasiperiodic system, which arise from generalized incommensurate zeros (GIZs) in matrix elements. This universal mechanism opens up a much broader way of rigorously realizing and characterizing the exotic critical states. Third, we establish the exactly solvable condition, under which the 1D spinful QP system can be transformed into effectively spinless QP models of dressed particles, and then can be exactly solved. The universal results provide a powerful guiding principle to construct exactly solvable models for novel localization physics, particularly the spin-selective QP model and QP optical Raman lattice model encompassing all basic types of MEs and all seven fundamental localization phases, respectively.
@@ II.1 The spinful quasiperiodic model
$H=\sum_{j,s,s^{\prime}}\big{(}c_{j+1,s}^{\dagger}\Pi_{j}^{s,s^{\prime}}c_{j,s^{\prime}}+\mathrm{h.c.}\big{)}+\sum_{j,s,s^{\prime}}c_{j,s}^{\dagger}M_{j}^{s,s^{\prime}}c_{j,s^{\prime}},$ (1)
where $c_{j,s}^{\dagger}(c_{j,s})$ creates (annihilates) a particle on site $j$ with (pseudo)spin $s=\uparrow,\downarrow$ that can denote real spin, sublattices or particle-hole. This model includes two key ingredients, as illustrated in Fig. 1(a). First, the hopping coupling matrix $\Pi_{j}=\sum_{l}p_{j}^{l}\sigma_{l}$, as written in terms of Pauli matrices $\sigma_{l}$ with $l=\{0,x,y,z\}$, represents the spin-independent (for $l=0$) and spin-dependent (for $l=x,y,z$) hopping terms. The latter includes both spin-conserved ($l=z$) and spin-flip ($l=x,y$) terms, corresponding to the spin-orbit coupling in periodic systems. Second, the onsite matrix $M_{j}=\sum_{l}m_{j}^{l}\sigma_{l}$ also involves the spin-independent chemical potential and the Zeeman coupling terms. The coefficients $p_{j}^{l}=t^{l}+\mu^{l}V_{j}^{\mathrm{od}}$ and $m_{j}^{l}=\lambda^{l}+V^{l}V_{j}^{\mathrm{d}}$, where $(t^{l},\lambda^{l},\mu^{l},V^{l})$ are constants, the off-diagonal QP modulation $V_{j}^{\mathrm{od}}=\cos[2\pi\alpha(j+1/2)+k_{y}]$ and the diagonal QP modulation $V_{j}^{\mathrm{d}}=\cos(2\pi\alpha j+k_{y})$. Here $\alpha=\lim_{n\rightarrow\infty}(F_{n-1}/F_{n})=(\sqrt{5}-1)/2$ is an irrational number approached by the Fibonacci sequence $F_{n}$, and $k_{y}$ is a constant phase factor. In this work, we take that the QP term has a single frequency. To facilitate our discussions, we define the combinations of Pauli matrices $\sigma_{l}$ as $\sigma_{\pm}=(\sigma_{x}\pm+i\sigma_{y})/2$ and $\Lambda_{\pm}=(\sigma_{z}\pm\sigma_{0})/2$.
This spinful QP system unifies various QP models, including the existing 1D spinful models with analytical solutions, and spinless models by taking the Majorana representation (see the details in Appendix A and Appendix F). More importantly, with this unified framework we shall develop a series of new and exactly solvable models which host various novel quantum phases with or without mobility edges. Below, we outline the three main theoretical approaches employed in this study.
@@ II.2 Dual transformation
The dual transformation is a non-local operation similar to the Fourier transformation, and maps the system between real space and dual momentum space. It is defined as $c_{j,s}=(1/\sqrt{L})\sum_{n}e^{-i2\pi n\alpha j}c_{n,s}$. The extended, localized, and critical states behave differently under this transformation. For an extended state, the wave function is delocalized in real space, but localized in dual momentum space, whereas the localized state is delocalized in dual momentum space but localized in real space. These opposing features are exchanged under the dual transformation. In contrast, critical states exhibit behavior distinct from the other two; they remain delocalized in both real and dual spaces, and are invariant under the transformation. For example, when the system has a QP hopping term $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$, which dominates over other hopping couplings and is invariant under the dual transformation due to the presence of incommensurately distributed zeros (IDZs) in the QP hopping modulation $V_{j}^{\rm od}$, the critical states may emerge and exhibit self-dual for having incommensurate nodes corresponding to the IDZs Zhou et al. (2023); Liu (2024). These incommensurate nodes in the delocalized wave functions remain unchanged under the dual transformation, driving the delocalized states into critical states. An intuitive picture suggests that IDZs partition the 1D system into multiple segments, where the wave function reorganizes, exhibiting a self-similar pattern characteristic of the critical states.
@@
The distinct behaviors of the three types of states can be mapped onto a 2D system to get an intuitive physical interpretation, where $x$- and $y$-axes correspond to the real and dual spaces, and the dual transformation exchanges the two axes. In this mapping, the extended state is delocalized along $x$-direction and localized along $y$-direction, and the opposite is true for localized states. The critical state is delocalized in both directions. Thus, the swapping of the two axes transforms the localized and extended states, while leaving the critical states unchanged. We use 1D spinless 1D extended Aubry-André (EAA) model Hatsugai and Kohmoto (1990); Han et al. (1994); Liu et al. (2015) as an example, which is characterized by the nearest-neighbor hopping $t_{j}=t+\mu V_{j}^{\rm od}$ and on-site potential $V_{j}=V_{0}V_{j}^{\rm d}$ (see details in Appendix A). By interpreting the phase shift $k_{y}$ as the quasi-momentum in the $y$-direction and performing a Fourier transformation, the 1D spinless EAA model maps to a 2D Hofstadter model that describes a particle moving in a magnetic field (Kraus and Zilberberg, 2012; Borgnia and Slager, 2023). In this mapping, the uniform hopping terms correspond to hopping in the $x$-direction, while the on-site potential $V_{j}^{\mathrm{d}}a_{j}^{\dagger}a_{j}$ corresponds to the hopping in the $y$-direction with a magnetic flux $\exp(i2\pi\alpha x)a_{x,y}^{\dagger}a_{x,y+1}+{\rm h.c.}$. Similarly, the QP hopping term $(V_{j}^{\mathrm{od}}a_{j}^{\dagger}a_{j+1}+\mathrm{h.c.})$ represents diagonal hopping in the 2D plane under the magnetic flux. The relative hopping amplitude shapes the cyclotron orbits while maintaining the flux within each orbit. When the hopping in the $x$- or $y$- direction dominates, the extended or localized states emerge. In contrast, when the hopping in diagonal direction dominates, hence having equal weight in both directions, the critical states are resulted and exhibit delocalized feature in both axes.